# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision Language Models

### v2

@@ Abstract
Large Vision Language Models (LVLMs), despite their recent success, are hardly comprehensively tested for their cognitive abilities. Inspired by the prevalent use of the “Cookie Theft” task in human cognition test, we propose a novel evaluation benchmark to evaluate high-level cognitive ability of LVLMs using images with rich semantics 111 Code and data are available at
https://github.com/xiujiesong/Cog-Bench.
. It defines eight reasoning capabilities and consists of an image description task and a visual question answering task. Our evaluation on well-known LVLMs shows that there is still a large gap in cognitive ability between LVLMs and humans.
A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision Language Models
Xiujie Song1, Mengyue Wu2thanks: Corresponding Authors., Kenny Q. Zhu311footnotemark: 1, Chunhao Zhang4, Yanyi Chen5 1,2,4Shanghai Jiao Tong University, Shanghai, China 3University of Texas at Arlington, Arlington, Texas, USA 5Rutgers University, New Brunswick, New Jersey, USA {1xiujiesong, 2mengyuewu, 4forest_zch}@sjtu.edu.cn 3kenny.zhu@uta.edu 5yanyi.chen@rutgers.edu
@@ 1 Introduction
Recently, with the emergence of Large Language Models (LLMs) such as GPT-4 OpenAI (2023), the cognitive abilities of language models have reached a new level Zhuang et al. (2023). They demonstrated remarkable performance in many tasks Bubeck et al. (2023). In Vision Language (VL), several researchers Zhu et al. (2023b); Liu et al. (2023b); Ye et al. (2023) endeavor to boost Vision Language Pre-trained Models (VLPMs) by integrating powerful LLMs Touvron et al. (2023); Chiang et al. (2023), referred to as Large Vision-Language Models (LVLMs). With LLM as the “brain”, the cognitive abilities of LVLMs are also improved and more challenging tasks, such as table, chart, and document reasoning etc, can be solved Yang et al. (2023). Some state-of-the-art LVLMs, such as GPT-4V OpenAI (2023), are progressing towards human-level cognitive abilities. Thus, there’s an increasing interset in evaluating the cognitive abilities of LVLMs. Though some LVLM evaluation benchmarks, such as MME Fu et al. (2023), MMBench Liu et al. (2023c), SEED Bench Li et al. (2023a), etc., also consider cognitive reasoning ability as one aspect of their evaluation, they do not provide a comprehensive evaluation of higher-level reasoning ability via a concentrated task and most of the images they use require relatively little reasoning to understand. However with humans, we have such tasks like Cookie Theft picture description, which is broadly used to assess one’s cognitive function.
Among different picture description tasks, the Cookie Theft picture description task (Figure 3) from the Boston Diagnostic Aphasia Examination Goodglass et al. (2001) has dominated work in speech-language pathology Cummings (2019). Surprisingly, this picture has been developed and utilized for more than 50 years and continues to be applied and discussed today. The success of Cookie Theft has sparked our curiosity to investigate the magic behind this picture and explore ways to apply this human-oriented cognitive ability assessment to evaluate cognitive abilities of LVLMs.
@@
Cummings (2019) analyzed the reasons for the success of of Cookie Theft. It was found that it contains information of varying degrees of salience and contains more semantic categories. It was also found that during describing Cookie Theft, compared to individuals with cognitive impairments, healthy people tend to demonstrate their coginitive ability by reasoning. For instance, in Figure 3, by comparing the descriptions produced by the healthy man and the woman with probable Alzheimer’s dementia, we can identify the following differences:
@@
Through these reasoning processes, the difference of cognitive abilities between the two individuals is reflected in their descriptions.
Such a picture that can evaluate cognitive function is carefully designed and crafted. Tasnim et al. (2022) introduced guidelines for drawing pictures similar to Cookie Theft, which is consistent with findings mentioned above. Generally speaking, compared to typical images, Cookie Theft-like images feature a prominent story theme, richer content, display relationships between components and require stronger cognitive abilities to understand and describe.
The above studies provided us with directions for collecting Cookie Theft-like images and developing assessment criteria to construct a cognitive evaluation benchmark, so that the idea of this human cognition assessment can be leveraged to evaluate cognition of LVLMs comprehensively. To fill this research gap, we propose to construct a Cognitive Evaluation Benchmark, named as CogBench, which consists of high quality Cookie Theft-like images to evaluate the cognitive reasoning ability of LVLMs. CogBench defines eight core cognitive reasoning capabilities, including reasoning about special time, location, character, character relationship, event, event relationship, next moment event and mental state. Both a generative Image Description task and a discriminative Visual Question Answering (VQA) task are designed.
Our main contributions are as follows:
  • To the best of our knowledge, this is the first-of-its-kind attempt to incorporate the cognitive test Cookie-Theft picture description task designed for humans into the evaluation of LVLMs.
  • We created the first ever and the largest to-date VL dataset with Cookie Theft-like images to evaluate LVLM’s cognitive abilities.
  • Our evaluation on existing well-known LVLMs shows that there is still a large gap between the cognitive ability of LVLMs and human beings, indicating CogBench is a challenging evaluation benchmark.
@@ 2 Dataset Construction
In this section, we will introduce the construction of CogBench. We will first introduce image collection, annotation and the statistics of CogBench. Then, we will introduce tasks in CogBench.
@@ 2.1 Image Collection
Based on findings of previous studies Cummings (2019); Tasnim et al. (2022), we set the following image collection criteria:
  • Rule 1: Stroy-telling The image depicts an interesting story clearly. For instance, the Cookie Theft picture tells the story of a mother busy washing dishes while two kids takes the opportunity to stand on a stool and sneakily steal cookies.
  • Rule 2: Rich Chain-of-Reasonings Images should display rich Chain-of-Reasonings (CoRs) in a scene. A CoR connects low-level observations in an image to produce a high-level reasoning conclusion or connects the cause and effect of events. For example, “The mother is busy washing dishes. + The boy is standing on the stool behind the mother. + The girl standing by the boy is shushing him. + The boy is fetching cookies from the jar in the carbinet. $\rightarrow$ The boy and girl are stealing cookies.” is a CoR about high-level event “stealing cookies”. Note that the story is actually constructed via these CoRs.
  • Rule 3: Content Complexity Restriction Images should contain rich content but not be overly complex. The number of subjects should be limited for better emphasis on the key points of the story, facilitating description by humans or models.
With the above criteria, we aim to select high-quality images for cognition evaluation of LVLMs. Most of the images are manully collected from Pinterest [^4] https://www.pinterest.com/, and the Cookie Theft is also included in CogBench. Figure 2 illustrates the differences between our images and those from other datasets.
@@ 2.2 Image Annotation
Human annotators, mostly undergraduate or graduate students aged 18-28, are hired to annotate the collected images. As shown in Figure 3, the annotation includes three parts: [Entities], [CoRs] and [Description].
By annotating [Entities] and [CoRs], we aim to evaluate the low-level recognition ability and high-level cognitive reasoning ability of models respectively based on description. [Description] is annotated as the reference description for the image. The three parts are annotated in a sequential order.
Entity Annotation
We ask annotators to list as many entities in the image as possible and entities that are difficult to recognize should be omitted.
CoR Annotation
In order to evaluate model cognition in a fine-grained manner, the following eight reasoning capabilities are annotated with CoRs:
  • Special Time Reasoning: reasoning about the special time of the story in the image, e.g. festivals, seasons etc.
  • Location Reasoning: reasoning about the location of the story in the image, e.g. near the school etc.
  • Character Reasoning: reasoning about the character of the subjects in the image, e.g. police officer etc.
  • Character Relationship Reasoning: reasoning about relationship between characters in the image, e.g. “the woman is the mother of the kids.”
  • Event Reasoning: reasoning about high-level events happened in the current moment and previous moments in the image based on clues in the picture. The difference between high-level and low-level lies in how much semantic information the event contains, e.g. “stealing cookies” is a higher-level event compared to “taking cookies” as it additionally conveys the semantic of “taking advantage without permission or knowledge.”
@@
  • Mental State Reasoning: reasoning about mental state of subjects in the image, e.g. daydreaming etc.
Description Summary
Annotators are last asked to write a description to tell the story in the image based on the annotated [Entities] and [CoRs]. The annotation instruction for annotators is shown in Appendix A.
Considering different people may have different understanding about some images, we ask three annotators to annotate each image. Then, we manually merge the three annotations as one final annotation mainly based on the principle of majority vote. For the [Description], if two or more annotators understand the story in the image in the same way, we accept one of the best descriptions as the final description. For [Entities] and [CoRs], we first accept most of the entities and CoRs that are annotated by at least two annotators. Other entities and CoRs are also included if they are reasonable.
@@ 2.3 CogBench Statistics
To build CogBench, we have gathered 95 images with 1041 entities, 881 CoRs, 95 descriptions and 1091 questions. Table 1 shows the distribution of these CoRs and questions. Though the number of images in CogBench is not large, the content contained in each picture is very rich. The number of CoRs in event-related reasoning and [Mental State Reasoning] is large, which is a manifestation of the rich interesting stories in the images.
@@
CoR 23 79 42 102 245 172 56 162
QA 47 94 70 159 258 168 74 221
Table 1: Distribution of CoRs and questions in CogBench.
2.4 Tasks in CogBench
To comprehensively evaluate cognitive abilities of LVLMs, we designed a generative Image Description task and a discriminative Multiple-Choice Question Answering task in CogBench.
2.4.1 Image Description Task
Inspired by Cookie Theft picture description task, we propose to assess the cognitive abilities of LVLMs through an Image Description task using Cookie Theft-like images. Recently, some researchers Xie et al. (2022); Zhu et al. (2023a); Zhuge et al. (2023) are also trying to improve model performance on image description task. The difference between our description task and existing image description task is that we expect LVLMs understand and describe the story in the image through high-level cognitive reasoning. For instance, in Figure 3, the description of the image should not only include what is in the picture but also focus on elucidating the story of “parents are going out to attend a formal event, and the mother is refusing the boy to accompany them” through a series of reasonings.
2.4.2 Visual Question Answering Task
All questions in VQA task are Multiple-Choice Questions with four options, which simplifies evaluation. Similar to Description task, the questions in VQA task are also related to high-level cognitive reasoning. For example, in Figure 4, we can ask question about the event “What is the woman doing?” or question about the cause of the event “Why is the mom placing her hands on the boy’s forehead?”, which require models to answer with reasoning.
Figure 4: An example of VQA task from CogBench.
We use a semi-automatic GPT-assisted approach to generate questions for images based on [CoRs] we annotated in the previous section, referred to as CoR-based GPT-assisted Question Generation. The main idea of the question generation is that for each CoR, we can ask questions about both the conclusion (right part of $\rightarrow$) and reasons to draw the conclusion (left part of $\rightarrow$), as shown in Fiure 4. The process can be divided into two stages.
In the first stage, we use GPT-4 OpenAI (2023) to generate questions for images in CogBench. For each category of reasoning capability, we design a sepcific prompt for GPT-4 to generate questions, so that questions related to different types of CoRs can be generated. For Question Generation, the key point is to generate high quality distractors. Thus, we encourage GPT-4 to hallucinate to generate more perplexing distractors in our prompt. Appendix B shows an example of prompt we use for GPT-4 to generate questions.
Though GPT-4 is powerful, it is still difficult for it to generate 100% satisfying questions for all images. Thus, in the second stage, we manually select and modify the questions generated by GPT-4. The principle of selection and modification is that the question cannot apparently point to the correct option and distractors should be as closely related to the question and the correct option as possible, thereby possessing a certain level of misleadingness. During selection, ChatGPT is also utilized to help detect those simple questions that can be easily answered without having to accept the image as input.
3 Evaluation Strategy of CogBench
In this section, we will introduce the evaluation strategy of tasks in CogBench.
3.1 Evaluation for Description Task
For the evluation of Description task, we consider model performance comprehensively on two levels: low-level Recognition ability and high-level Cognition ability. Evaluation metrics for both levels are calculated based on recall score, referred to as Recognition Score and Cognition Score, respectively.
The Recognition Score is calculated as the ratio of the number of recognized [Entities] to the number of annotated [Entities] in all images.
First, we use SpaCy [^5] https://spacy.io/ to extract nouns from the model generated description, and then calculate cosine similariy between embeddings [^6] Implemented with sentence-transformers package (https://www.sbert.net) and all-mpnet-base-v2 is adopted as the model to encode [Entities] and nouns. of annotated [Entities] and extracted nouns. For each entity, if the cosine similarity score between the entity and any noun is greater than a threshold (0.5 in this paper), we consider the entity is recognized by the model.
For Cognition Score, we first calculate a score for each of the eight cognitive reasoning abilities, and then compute an overall score.
We use GPT-4 OpenAI (2023) to help calculate the Cognition Score. To avoid the interpretability issues of GPT-based evaluation, instead of using a subjective evaluation method of directly comparing two descriptions, we use GPT-4 to perform a easier, more objective and fine-grained binary classification task, that is, judging whether the generated description contains the semantics of annotated CoRs. Details and prompts of this GPT-based evaluation are shown in Appendix C.
After obtaining the score of each CoR, we calculate a recall score for each reasoning capabilities. The overall Cognition Score is calculated by summing up the scores of all CoRs and dividing by the total number of CoRs. The effectiveness analysis of GPT-based cognition evaluation is shown in Appendix F.
3.2 Evaluation for VQA Task
For Multiple-Choice Questions in VQA task, we use accuracy as the evaluation metric. As questions are generated based on CoRs, we can also calculate both the accuracy for each reasoning capability and the overall accuracy.
@@ 4 Experiments
In this section, we will introduce selected LVLMs for evaluation, experiment results and analysis.
4.1 Large Vision Language Models
InstructBLIP-7B Dai et al. (2023), Qwen-VL-Chat Bai et al. (2023), LLaVA-v1.5-7B and LLaVA-v1.5-13B Liu et al. (2023a), mPLUG-Owl-2 Ye et al. (2023), ShareGPT4V-7B and ShareGPT4V-13B Chen et al. (2023a) and GPT-4V OpenAI (2023) are selected in this paper to be evaluated using CogBench. A brief introduction of these models is shown in Appendix D.
@@ 4.2 Results of Description Task
We prompt the selected LVLMs with the following instruction to obtain detailed descriptions about images in CogBench.
Describe this image in detail.
Then, we evaluate the performance of LVLMs on Description task in terms of both recognition and cognition ability.
As a reference, we also calculated traditional image captioning evaluation metrics by camparing annotated reference description and model-generated description, and details are shown in Appendix E.

### v3

# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models

@@ Abstract
Large Vision-Language Models (LVLMs), despite their recent success, are hardly comprehensively tested for their cognitive abilities. Inspired by the prevalent use of the “Cookie Theft” task in human cognition test, we propose a novel evaluation benchmark to evaluate high-level cognitive ability of LVLMs using images with rich semantics 111 The code and the data of CogBench is released at: https://github.com/xiujiesong/Cog-Bench. . It defines eight reasoning capabilities and consists of an image description task and a visual question answering task. Our evaluation on well-known LVLMs shows that there is still a large gap in cognitive ability between LVLMs and humans.
@@ 1 Introduction
Recently, with the emergence of Large Language Models (LLMs) such as GPT-4 [1], the cognitive abilities of language models have reached a new level [2]. They demonstrated remarkable performance in many tasks [3]. In Vision Language (VL), several researchers [4, 5, 6] endeavor to boost Vision-Language Pre-trained Models (VLPMs) by integrating powerful LLMs [7, 8], referred to as Large Vision-Language Models (LVLMs). With LLM as the “brain”, the cognitive abilities of LVLMs are also improved and more challenging tasks, such as table, chart, and document reasoning, can be solved [9]. Some state-of-the-art LVLMs, such as GPT-4V and GPT-4o[1], are progressing towards human-level cognitive abilities [9]. Thus, there’s a growing interest in evaluating the cognitive abilities of LVLMs. Such evaluation usually amounts to co-reasoning of images and text, e.g., question answering about a picture. Though some LVLM evaluation benchmarks, such as MME [10], MMBench [11], and SEED Bench [12], evaluate cognitive reasoning ability as one aspect of their evaluation, they do not provide a comprehensive evaluation of higher-level reasoning abilities and most of the images they use require relatively little reasoning to understand.
In our study, we draw inspiration from the “Cookie Theft” picture description task (Figure 3), a key component of the Boston Diagnostic Aphasia Examination [13], widely utilized in clinical practice within speech-language pathology [14] and cognitive function screening [15]. Notably, despite being designed over half a century ago, this picture remains prevalent in contemporary psychological discussions.
@@
Is it possible to transfer the success of Cookie Theft picture description in human cognition test to evaluating cognitive abilities of LVLMs? Linguists and psychologists [14] conducted an analysis to determine the factors contributing to the success of “Cookie Theft”. The study revealed that the narrative includes information with varying levels of importance and encompasses a broad range of semantic categories. It was observed that during the description of “Cookie Theft”, individuals with intact cognitive functions exhibit their cognitive prowess by logically deducing the events and their interconnections. In contrast, those with cognitive impairments tend to merely list the superficial aspects of the situation without deeper reasoning. For instance, in Figure 3, by comparing the descriptions produced by the healthy man and the woman with probable Alzheimer’s dementia, we can identify the following differences:
@@
Through these reasoning processes, the difference of cognitive abilities between the two individuals is reflected in their descriptions. A picture that can evaluate cognitive functions needs to be carefully designed and crafted. Tasnim et al. [16] introduced guidelines for drawing pictures similar to Cookie Theft, which is consistent with findings mentioned above. Generally speaking, compared to ordinary images, Cookie Theft-like images feature i) a prominent story theme, ii) richer content, iii) display complex relationships among entities, and thus require stronger cognitive abilities to understand and describe.
With the above design principles, we propose to construct a Cognitive Evaluation Benchmark, named as CogBench, which consists of high quality Cookie Theft-like images to evaluate the cognitive reasoning ability of LVLMs. CogBench defines eight core cognitive reasoning capabilities, including reasoning about special time, location, character, character relationship, event, event relationship, next moment event and mental state. Both a generative Image Description task and a discriminative Visual Question Answering (VQA) task are designed. Our main contributions are as follows:
  • To the best of our knowledge, this is the first-of-its-kind attempt to incorporate a cognitive test designed for humans into the evaluation of LVLMs.
  • We created and open-sourced the largest VL dataset with Cookie Theft-like images for evaluating LVLM’s cognitive abilities.
  • Our evaluation on existing LVLMs shows that a large gap exists between the cognitive ability of LVLMs and human beings, indicating CogBench will be a valuable evaluation benchmark in the near future.
@@ 2 Dataset Construction
In this section, we will introduce the construction of CogBench. We will introduce image collection, annotation, tasks and the statistics of CogBench. Ethical considerations of CogBench are discussed in Appendix G.
@@ 2.1 Image Collection
Based on previous studies [14, 16], we set the following image collection criteria:
  • Rule 1: Story-telling The image depicts an interesting story. For instance, the Cookie Theft picture tells the story of a mother busy washing dishes while two kids takes the opportunity to stand on a stool and sneakily steal cookies.
  • Rule 2: Rich Chain-of-Reasonings Images should display rich Chain-of-Reasonings (CoRs) in a scene. A CoR connects low-level observations in an image to produce a high-level reasoning conclusion or connects the cause and effect of events. For example, “The mother is busy washing dishes. + The boy is standing on the stool behind the mother. + The girl standing by the boy is shushing him. + The boy is fetching cookies from the jar in the cabinet. $\rightarrow$ The boy and girl are stealing cookies.” is a CoR about high-level event “stealing cookies”. Note that the story is actually constructed via these CoRs.
  • Rule 3: Content Simplicity Images should contain rich content but not be overly complex. The number of subjects should be limited for better emphasis on the key points of the story.
With the above criteria, we manually collect pictures from Pinterest [^4] https://www.pinterest.com/, and the Cookie Theft is also included in CogBench. Figure 2 shows the differences between our images and those from other datasets by examples.
@@ 2.2 Image Annotation
Human annotators, mostly undergraduate or graduate students aged 18-28, are hired to annotate the collected images [^5] Wages and recruitment process details can be found in Appendix G.. As shown in Figure 3, the annotation includes three parts: [Entities], [CoRs] and [Description]. By annotating [Entities] and [CoRs], we aim to evaluate the low-level recognition ability and high-level cognitive reasoning ability of models respectively based on description. [Description] is annotated as the reference description for the image. The three parts are annotated in that order.
Entity Annotation We ask annotators to list as many entities in the image as possible and entities that are difficult to recognize should be omitted.
CoR Annotation In order to evaluate model cognition in a fine-grained manner, the following eight reasoning capabilities are annotated with CoRs:
  • Special Time Reasoning: reasoning about the special time of the story in the image, e.g., festivals, seasons etc.
  • Location Reasoning: reasoning about the location of the story in the image, e.g., near the school etc.
  • Character Reasoning: reasoning about the character of the subjects in the image, e.g., police officer etc.
  • Character Relationship Reasoning: reasoning about relationship between characters in the image, e.g., “the woman is the mother of the kids.”
  • Event Reasoning: reasoning about high-level events in the current moment and previous moments in the image based on clues in the picture. The difference between high-level and low-level lies in how much semantic information the event contains, e.g., “stealing cookies” is a higher-level event compared to “taking cookies” as it additionally conveys the semantic of “taking advantage without permission or knowledge.”
@@
  • Mental State Reasoning: reasoning about mental state of subjects in the image, e.g., daydreaming etc.
Description Summary Annotators are last asked to write a description to tell the entire story in the image based on the annotated [Entities] and [CoRs].
The complete annotation instruction is in Appendix A. Considering different people may have different understanding about some images, we ask three annotators to annotate each image. Then, we manually merge the three annotations as one final annotation by majority vote. For the [Description], if two or more annotators understand the image in the same way, we accept the best descriptions as the final description. For [Entities] and [CoRs], we first accept most of the entities and CoRs that are annotated by at least two annotators. Other entities and CoRs are also included if they are reasonable. For images with significantly different understandings from the three annotators, we will discard these images.
2.3 Tasks in CogBench
We design a generative Image Description task and a discriminative Multiple-Choice Question Answering task in CogBench.
2.3.1 Image Description Task
This is the primary task of the benchmark. The difference between our description task and existing image description task [17, 18, 19] is that we expect LVLMs understand and describe the story in the image through high-level cognitive reasoning. For instance, in Figure 3, the description of the image should not only include what is in the picture but also focus on elucidating the story of “parents are going out to attend a formal event, and the mother is refusing the boy to accompany them” through a series of reasonings.
2.3.2 Visual Question Answering Task
Figure 4: Generating a MCQ from event reasoning annotations.
The VQA task features standard four-option Multiple-Choice Questions, easing the evaluation process. Like the Description task, VQA questions involve different types of high-level cognitive reasoning, as illustrated by the question about events in Figure 4.
[…]
@@ 2.4 CogBench Statistics
Table 1: Distribution of CoRs and questions in CogBench.
@@
CoR 47 179 106 263 701 425 107 417
QA 86 220 162 317 658 402 135 597
CogBench consists of 251 semantically-rich images with a total of 2670 entities, 2245 CoRs, 251 descriptions and 2577 questions, indicating the content contained in each image is complex, showcased in Table 1. The number of CoRs in event-related reasoning and [Mental State Reasoning] is large, which is a manifestation of the rich interesting stories in the images.
@@ 3 Experiments
In this section, we will evaluate selected LVLMs, detail evaluation metrics and discuss the results.
3.1 Large Vision-Language Models
We evaluated a selection of recent representative open-source and closed-source LVLMs in this paper, including InstructBLIP-7B [20], Qwen-VL-Chat [21], LLaVA-v1.5-7B and LLaVA-v1.5-13B [22], mPLUG-Owl-2 [6], ShareGPT4V-7B and ShareGPT4V-13B [23], GPT-4V and GPT-4o [1]. A brief introduction of these models is in Appendix D. These open-source models are deployed on a machine with four GeForce RTX 2080 Ti GPUs and a machine with one GeForce RTX 3090 GPU.
3.2 CogBench Evaluation Strategy
Evaluation for Description Task
Here we consider model performance at two levels: low-level Recognition ability and high-level Cognition ability. Evaluation metrics for both levels are calculated using recall scores, referred to as Recognition Score and Cognition Score, respectively.
The Recognition Score is calculated as the ratio of the number of recognized [Entities] to the number of annotated [Entities] in all images. First, we use SpaCy [^6] https://spacy.io/ to extract nouns from the model generated description, and then calculate cosine similarity between embeddings [^7] Implemented with sentence-transformers package (https://www.sbert.net) and all-mpnet-base-v2 is adopted as the model to encode [Entities] and nouns. of annotated [Entities] and extracted nouns. For each entity, if the cosine similarity score between the entity and any noun is greater than a threshold (0.6 in this paper), we consider the entity is recognized by the model.
For Cognition Score, we calculate both scores for each of the eight cognitive reasoning abilities and an overall score using GPT-4 [1]. To enhance objectivity and granularity, GPT-4 is utilized for a binary classification task to assess if generated descriptions include the semantics of annotated CoRs, as detailed in Appendix C. The CoR scores for each type of reasoning capability are then used to compute a recall score for each respective type. The overall Cognition Score is the sum of all CoR scores divided by the total number of CoRs.
Evaluation for VQA Task
For Multiple-Choice Questions in VQA task, we use accuracy as the evaluation metric. As questions are generated based on CoRs, we can both calculate the accuracy for each reasoning capability and the overall cognition capability.
@@ 3.3 Results of Description Task
We prompt the selected LVLMs with the following instruction to obtain detailed descriptions about images in CogBench: ‘‘Describe this image in detail.’’ We evaluate the performance of LVLMs on Description task in terms of both recognition and cognition ability. As a reference, we also calculated traditional image captioning evaluation metrics by comparing annotated reference description and model-generated description, and details are shown in Appendix E.